# The Exact Relativistic Scalar Quasibound States of The Dyonic Kerr-Sen Black Hole: Quantized Energy, and Hawking Radiation

### v1

@@ 4. For the case with $\alpha_{+},\beta_{-},\gamma_{-}$, we obtain,
$\Omega_{2}=\frac{r_{s}\left(-in(r_{+}-r_{-})+2m_{l}a\right)}{2\left(r_{-}(r_{-}-4d)+2r_{+}d+a^{2}-k^{2}\right)}.$ (124)
@@ 5. For the case with $\alpha_{-},\beta_{+},\gamma_{+}$, we obtain,
$\Omega_{2}=\frac{r_{s}\left(in(r_{+}-r_{-})+2m_{l}a\right)}{2\left(r_{-}(r_{-}-4d)+2r_{+}d+a^{2}-k^{2}\right)}.$ (126)
@@ 8. For the case with $\alpha_{-},\beta_{-},\gamma_{-}$, we obtain,
$\Omega_{2}=\frac{r_{s}\left(-in(r_{+}-r_{-})+2m_{l}a\right)}{2\left(r_{-}(r_{-}-4d)+2r_{+}d+a^{2}-k^{2}\right)}.$ (130)
@@ IV The Horizon’s Hawking Radiation
$R=\left\{\begin{array}[]{cc}\psi_{+in}=A{\left(\frac{r_{+}-r_{-}}{\delta_{r}}\right)}^{\frac{1}{2}\gamma}{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{\frac{1}{2}\beta}&\emph{ingoing}\\ \psi_{+out}=B{\left(\frac{r_{+}-r_{-}}{\delta_{r}}\right)}^{\frac{1}{2}\gamma}{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{-\frac{1}{2}\beta}&\emph{outgoing}\end{array}\right..$ (136)
@@
$\begin{split}{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{\lambda}&={\left(\frac{r_{+}}{\delta_{r}}\right)}^{\lambda}{\left(\frac{r}{r_{+}}-1\right)}^{\lambda}\\ &\to{{\left(\frac{r_{+}}{\delta_{r}}\right)}^{\lambda}\left[\left(\frac{r}{r_{+}}-1\right)+i\epsilon\right]}^{\lambda}\\ &=\left\{\begin{array}[]{cc}{\left(\frac{r_{+}}{\delta_{r}}\right)}^{\lambda}{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{\lambda}&,\ r>r_{+}\\ {\left(\frac{r_{+}}{\delta_{r}}\right)}^{\lambda}{\left\lvert\frac{r-r_{+}}{\delta_{r}}\right\rvert}^{\lambda}e^{i\lambda\pi}&,\ r<r_{+}\end{array}\right.,\end{split}$ (137)
The analytical continuation enables us to obtain the expression $\psi_{-out}=\psi_{+out}\left(\left(\frac{r-r_{+}}{\delta_{r}}\right)\to\left(\frac{r-r_{+}}{\delta_{r}}\right)e^{i\pi}\right)$ simply by replacing $\left(\frac{r-r_{+}}{\delta_{r}}\right)\to-\left(\frac{r-r_{+}}{\delta_{r}}\right)=\left(\frac{r-r_{+}}{\delta_{r}}\right)e^{i\pi}$ as follows,
$\begin{split}\psi_{-out}&=B{\left(\frac{r_{+}-r_{-}}{\delta_{r}}\right)}^{\frac{1}{2}\gamma}{\left(\left(\frac{r-r_{+}}{\delta_{r}}\right)e^{i\pi}\right)}^{-\frac{1}{2}\beta},\\ &=\psi_{+out}e^{-\frac{1}{2}i\pi\beta}.\end{split}$ (138)
@@
Interestingly enough, similar to the case of the Kerr black hole, the angular momentum plays the role of the chemical potential in the thermodynamics of Kerr-Sen black hole, with $\mu=\hbar\omega_{J}$. For the Kerr-Sen black hole, $\mu$ also depends on the axion and dilaton (or electric and magnetic) charges of the black hole.
@@ V Conclusions
by applying the small black hole limit $Er_{s}\to 0$ and taking as zero the spin $a$, and the charges of the Dyonic Kerr-Sen black holes, $Q,P,k,d$, in the equations (116) and (118). The $\frac{1}{n^{2}}$ Hydrogenic-atom-like energy expression was also obtained in the previous works Daniel ; Anal1 ; Anal2 ; Anal3 ; Anal10 ; David-Senjaya .
@@ Acknowledgments
DS acknowledges this work is supported by the Second Century Fund (C2F), Chulalongkorn University, Thailand. P.B. is supported in part by National Research Council of Thailand (NRCT) and Chulalongkorn University under Grant N42A660500. This research has received funding support from the NSRF via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation [grant number B41G670027].
@@ Appendix A The Confluent Heun Functions and S-Homotopic Transformation
$y=e^{\pm i\sqrt{A_{5}}}x^{\pm i\sqrt{A_{3}}}(x-1)^{\pm i\sqrt{A_{4}}}\left[A\operatorname{HeunC}\left(\alpha,\beta,\gamma,\delta,\eta,x\right)\right.\\ \left.+Bx^{-\beta}\operatorname{HeunC}\left(\alpha,-\beta,\gamma,\delta,\eta,x\right)\right],$ (176)

### v2

@@ 4. For the case with $\alpha_{+},\beta_{-},\gamma_{-}$, we obtain,
$\Omega_{2}=\frac{r_{s}\left(-in(r_{+}-r_{-})+2m_{l}a\right)}{2\left(r_{-}(r_{-}-2d)+a^{2}-k^{2}\right)}.$ (124)
@@ 5. For the case with $\alpha_{-},\beta_{+},\gamma_{+}$, we obtain,
$\Omega_{2}=\frac{r_{s}\left(in(r_{+}-r_{-})+2m_{l}a\right)}{2\left(r_{-}(r_{-}-2d)+a^{2}-k^{2}\right)}.$ (126)
@@ 8. For the case with $\alpha_{-},\beta_{-},\gamma_{-}$, we obtain,
$\Omega_{2}=\frac{r_{s}\left(-in(r_{+}-r_{-})+2m_{l}a\right)}{2\left(r_{-}(r_{-}-2d)+a^{2}-k^{2}\right)}.$ (130)
@@
III.5 Numerical Investigations of the Black Hole’s Parameters
To investigate how the spin and charges of the Dyonic Kerr-Sen black hole affect a scalar field, let us choose an appropriate mode to investigate, i.e., the mode $\alpha_{+},\beta_{-},\gamma_{-}$, which has $\text{Re}(\Omega)>0$ that represents particle state. We investigate the QNMs by plotting the exact energy expression for various spin and charges ranged from zero to near-extremal configurations ($r_{-}\sim r_{+}$).
In Fig. 1, we present the real and imaginary parts of the quasiresonance (quasinormal) frequencies of $n=1,2,...,7$ states for various combinations of sub-extremal spin and charges (in geometrical unit), where the scalar and black hole mass are set to be respectively $r_{s}=1,\Omega_{0}=0.1$. Notice that the higher the excitation, the closer $\text{Re}(\Omega)$ to $\Omega_{0}$, indicating weaker binding energy for higher excited states.
Figure 1: QNMs for various combinations of sub-extremal spin and charges.
In Fig. 2,3,4,5, visualization of quasiresonance frequencies for $n=1,2,\dots,5$ for varied scalar mass, electric charge, magnetic charge and angular momentum, respectively, are presented. For each investigation, one black hole’s parameter is varied from zero up to its near-extremal limit, while the others are kept constant. In all cases, we observe significant changes happen in the region where the Dyonic Kerr-Sen black hole becomes nearly extremal. Also notice that the states with smaller radial quantum number $n$, i.e., low excited states, are affected more than the states with large $n$, i.e., higher excited states.
In Fig. 6,7,8, we investigate how the QNM profile changes as the Dyonic Kerr-Sen black hole becomes nearly extremal. Firstly, we fix the black hole’s mass, spin and the electric charge and we plot the quasiresonance frequencies with respect to the magnetic charge approaching the extremal limit followed by varying the electric charge and angular momentum. We consistently observe significant changes in states with smaller radial quantum number $n$ as the black hole becomes nearly extremal. Also notice that very close to the extremal limit, we observe there are fundamental states having $\text{Re}(\Omega)$ larger than $\Omega_{0}$ (see Fig. 6,7,8.)
Figure 2: QNMs for varied scalar mass.
Figure 3: QNMs for varied electric charge.
Figure 4: QNMs for varied magnetic charge.
Figure 5: QNMs for varied angular momentum.
Figure 6: Near extremal QNMs for varied magnetic charge.
Figure 7: Near extremal QNMs for varied electric charge.
Figure 8: Near extremal QNMs for varied angular momentum.
III.6 Analytical Investigations of the Black Hole’s Parameters
In this subsection, we will investigate analytically, how the black hole’s angular momentum and charges affect the scalar field’s energy levels. The exact expression of the energy levels, in general, are complex valued quartic equations. It is very difficult to algebraically figure out how the black hole spin and charges affect the scalar’s energy levels. However, it is possible to analytically investigate the approximated energy expression (116) in comparison with the well-known Schwarzschild’s gravitational atom expression in Daniel ; Anal1 ; Anal2 ; Anal3 ; Anal10 ; David-Senjaya ; Huang as follows,
${\left(E-E_{0}\right)}_{Sch}=-\frac{E_{0}}{8}{\left[\frac{E_{0}r_{S}}{\hslash c}\right]}^{2}{\left(\frac{1}{n}\right)}^{2},$ (131)
Since the Schwarzschild black hole is the Dyonic Kerr-Sen black hole with vanishing charges and angular momentum, the comparison between the two may give some insights on how the black hole spin and charges affect the scalar field’s quantized energy.
Now, let us consider the approximated energy expression in equation (116). This particular mode is important and interesting due to its dependence on the black hole’s angular momentum. The expression can be rewritten in $x+iy$ form as follows,
${\left(E-E_{0}\right)}_{DKS}$ $\approx-\frac{E_{0}}{2}{\left[\frac{\frac{E_{0}r_{S}}{\hslash c}\left(\frac{r_{+}+r_{-}-2d}{r_{s}}\right)}{2\left(i\frac{2m_{l}a}{{\delta}_{r}}+n\right)}\right]}^{2}$
$\approx-\frac{E_{0}}{8}{\left[\frac{E_{0}r_{S}}{\hslash c}\left(\frac{r_{+}+r_{-}-2d}{r_{s}}\right)\right]}^{2}\times$
$\phantom{=}\left(\frac{n^{2}-{n_{a}}^{2}}{{\left(n^{2}+{n_{a}}^{2}\right)}^{2}}-i\frac{2n_{a}n}{{\left(n^{2}+{n_{a}}^{2}\right)}^{2}}\right),$ (132)
where $n_{a}=\frac{2m_{l}a}{{\delta}_{r}}$ and by using the relation (19), we obtain $r_{+}+r_{-}-2d=r_{s}$.
Evaluating the ratio between ${\left(E-E_{0}\right)}_{Sch}$ and ${\left(E-E_{0}\right)}_{DKS}$, we obtain the following relation,
$\frac{{\left(E-E_{0}\right)}_{DKS}}{{\left(E-E_{0}\right)}_{Sch}}$ $=\ n^{2}\left(\frac{n^{2}-{n_{a}}^{2}}{{\left(n^{2}+{n_{a}}^{2}\right)}^{2}}-i\frac{2n_{a}n}{{\left(n^{2}+{n_{a}}^{2}\right)}^{2}}\right)$
$=\frac{1}{{\left(1+{\left(\frac{n_{a}}{n}\right)}^{2}\right)}^{2}}\left[1-{\left(\frac{n_{a}}{n}\right)}^{2}-2\frac{n_{a}}{n}i\right]$
$={\mathfrak{R}}_{Re}+i{\mathfrak{R}}_{Im},$ (133)
where,
${\mathfrak{R}}_{Re}=\frac{1-{\left(\frac{n_{a}}{n}\right)}^{2}}{{\left(1+{\left(\frac{n_{a}}{n}\right)}^{2}\right)}^{2}}\ \ \ ,\ \ \ {\mathfrak{R}}_{Im}=-2\frac{\frac{n_{a}}{n}}{{\left(1+{\left(\frac{n_{a}}{n}\right)}^{2}\right)}^{2}}.$ (134)
The function ${\mathfrak{R}}_{Re}$ and ${\mathfrak{R}}_{Im}$ for various value of excitation numbers are shown in Fig. 9.
Figure 9: The ratio between scalar’s binding energy in the Dionic Kerr-Sen with the Schwarzschild black hole in $(E-E_{0})r_{s}\to 0$ limit.
The variable ${\delta}_{r}=r_{+}-r_{-}$ which measures the distance between the outer and inner horizons has the following explicit expression,
${\delta}_{r}=2\sqrt{{\left(\frac{r_{s}}{2}\right)}^{2}+d^{2}+k^{2}-\left(a^{2}+r^{2}_{D}\right)}.$ (135)
Notice that the black hole’s angular momentum together with electric and magnetic charges in $r^{2}_{D}$ are in opposite sign to the mass, dilaton and axion charges. This indicates that the presence of dilaton and axion charges shifts the inner horizon further inward and the outer horizon further outward resulted in larger separation between $r_{+}$ and $r_{-}$, while the presence of the spin affects the horizon separation in the opposite way. The parameter ${\delta}_{r}$ is then scaling the magnetic-spin interaction $n_{a}=\frac{2m_{l}a}{{\delta}_{r}}$, that larger ${\delta}_{r}$ reduces $n_{a}$ and vice versa, and the special case with $m_{l}=0$ or $a=0$ will effectively null the interaction term.
The ratio ${\frac{{\left(E-E_{0}\right)}_{DKS}}{{\left(E-E_{0}\right)}_{Sch}}}$ tells us that for a scalar field in a general Dyonic Kerr-Sen space-time having main quantum number $n$ and non-zero magnetic quantum number $m_{l}$, the real part of ${\left(E-E_{0}\right)}_{DKS}$ will always be smaller than the same mode in the Schwarzschild space-time. There will also be additional decay for modes with $m_{l}>0$ and instability for modes with $m_{l}<0$. Thus, the presence of the black hole’s angular momentum and all charges affect fundamental mode, i.e. $n=1$, more than modes with larger $n$ since,
${\mathop{\mathrm{lim}}_{n\to\infty}{\mathfrak{R}}_{Re}\ }=1,\ \ \ \ \ {\mathop{\mathrm{lim}}_{n\to\infty}{\mathfrak{R}}_{Im}\ }=0.$ (136)
For the massless scalar case given by (119)-(130), the presence of non-zero black hole’s spin gives rise to QBS with non-zero real part, i.e. solutions for underdamped modes with $\beta_{+},\gamma_{+}$ and $\beta_{-},\gamma_{-}$. while the rest are not affected by the black hole’s spin and charges and only depend only on the black hole’s mass since $r_{+}+r_{-}-2d=r_{s}$.
@@ IV The Horizon’s Hawking Radiation
$R=\left\{\begin{array}[]{cc}\psi_{+in}=A{\left(\frac{r_{+}-r_{-}}{\delta_{r}}\right)}^{\frac{1}{2}\gamma}{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{-\frac{1}{2}\beta}&\emph{ingoing}\\ \psi_{+out}=B{\left(\frac{r_{+}-r_{-}}{\delta_{r}}\right)}^{\frac{1}{2}\gamma}{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{\frac{1}{2}\beta}&\emph{outgoing}\end{array}\right..$ (142)
@@
$\begin{split}{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{\lambda}&={\left(\frac{r_{+}}{\delta_{r}}\right)}^{\lambda}{\left(\frac{r}{r_{+}}-1\right)}^{\lambda}\\ &\to{{\left(\frac{r_{+}}{\delta_{r}}\right)}^{\lambda}\left[\left(\frac{r}{r_{+}}-1\right)-i\epsilon\right]}^{\lambda}\\ &=\left\{\begin{array}[]{cc}{\left(\frac{r_{+}}{\delta_{r}}\right)}^{\lambda}{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{\lambda}&,\ r>r_{+}\\ {\left(\frac{r_{+}}{\delta_{r}}\right)}^{\lambda}{\left\lvert\frac{r-r_{+}}{\delta_{r}}\right\rvert}^{\lambda}e^{-i\lambda\pi}&,\ r<r_{+}\end{array}\right.,\end{split}$ (143)
The analytical continuation enables us to obtain the expression $\psi_{-out}=\psi_{+out}\left(\left(\frac{r-r_{+}}{\delta_{r}}\right)\to\left(\frac{r-r_{+}}{\delta_{r}}\right)e^{-i\pi}\right)$ simply by $\left(\frac{r-r_{+}}{\delta_{r}}\right)\to-\left(\frac{r-r_{+}}{\delta_{r}}\right)=\left(\frac{r-r_{+}}{\delta_{r}}\right)e^{-i\pi}$ as follows,
$\begin{split}\psi_{-out}&=B{\left(\frac{r_{+}-r_{-}}{\delta_{r}}\right)}^{\frac{1}{2}\gamma}{\left(\left(\frac{r-r_{+}}{\delta_{r}}\right)e^{-i\pi}\right)}^{\frac{1}{2}\beta},\\ &=\psi_{+out}e^{-\frac{1}{2}i\pi\beta}.\end{split}$ (144)
@@
Interestingly enough, similar to the case of the Kerr black hole, the angular momentum plays the role of the chemical potential in the thermodynamics of Kerr-Sen black hole, with $\mu=\hbar\omega_{J}$ Navarro . For the Kerr-Sen black hole, $\mu$ also depends on the axion and dilaton (or electric and magnetic) charges of the black hole.
@@ V Conclusions
by applying the small black hole limit $Er_{s}\to 0$ and taking as zero the spin $a$, and the charges of the Dyonic Kerr-Sen black holes, $Q,P,k,d$, in the equations (116) and (118). The $\frac{1}{n^{2}}$ Hydrogenic-atom-like energy expression was also obtained in the previous works Daniel ; Anal1 ; Anal2 ; Anal3 ; Anal10 ; David-Senjaya ; Huang .
We have performed further numerical and analytical investigations on how the Dyonic Kerr-Sen black hole’s parameters affect the energy levels of the scalar fields. We plot graphical visualizations of QBS with various configurations of scalar field-Dyonic Kerr-Sen black hole and find out that in general, the states with larger radial quantum number $n$ have $\text{Re}(\Omega)$ closer to $\Omega_{0}$, indicating weaker binding energy 1. Very close to the extremal limit, we observe there are states with small radial quantum number having $\text{Re}(\Omega)$ larger than $\Omega_{0}$ 6,7,8. In general, we find that states with smaller radial quantum number $n$ are more sensitive to the change of the black hole’s parameters, which in agreement with the analytical investigation that has been done by making use of the approximated energy expression (116).
@@
Analytical solutions, and results, are extremely useful in the study of the astrophysical and physical properties of the black holes, including the study of the dynamics and motion of the particles gravitating around them. They can also be successfully used to investigate the emissivity properties in various frequency ranges of the thin accretion disks that generally are present around black holes. The obtained exact solution of the scalar quasibound states of the Dyonic Kerr-Sen black hole can also help in discriminating this type of black hole with respect to standard general relativistic black holes, or other black hole solutions in modified gravity theories, and for obtaining observational constraints on the black hole parameters, and on mass, spin and the various charges of the black hole.
We have also presented a detailed investigation of the thermodynamic properties of the Dyonic Kerr-Sen black hole, by using the obtained exact analytical results. The Hawking temperature is one of the extremely important physical parameters of the black holes, representing an essential quantity that has very important theoretical implications. In the limiting case of the vanishing angular momentum and charges we recover the expression of the Hawking temperature of the general relativistic Schwarzschild black holes. As compared to the standard Schwarzschild or Kerr cases, the horizon temperature of the Dyonic Kerr-Sen black holes has a strong dependence on the electric and magnetic charges, a new physical feature that could in principle discriminate between the various types of black holes, and their physical properties. The study of the thermodynamical properties of the black hole was very much facilitated by the existence of the exact solution.
Bound systems play an important role in classical mechanics, being fundamental for the understanding of the dynamics of the Sun and of the Solar System planets. The study of the bound states of the quantum mechanical systems, like, for example, the hydrogen atom, opened new perspectives on the behavior of elementary particles. The discovery of the gravitational waves, together with the confirmation of the existence of black holes by using various astrophysical methods and physical methods, led to a significant increase in the interest for bound and scattering states, and for quasinormal excitations of the particles gravitating around compact massive objects. There are no real bound states for particle around black holes, since they will cross into the black hole via quantum effects, leading to the slow, or rapid decay of the matter waves. Hence, these decaying states are called quasibound states. Generally, for astrophysical black holes, the mass of the quasibound particles is extremely small, and thus it is very difficult to detect them by using accelerator experiments.
However, the situation is drastically different in an astrophysical environment, and the physical processes around black holes may lead to the possibility of testing the existence and properties of ultra light particles. If they have proper frequencies, the particles gravitating around a hole could create a large number of similar ones. The type of radiation corresponding to these processes (superradiance) is a non-thermal one, and it is different from the Hawking radiation. Due to the black hole properties, some particles will cross the event horizon into the black hole by quantum tunnelling. However, if the two processes of creation and absorption are in detailed balance, a cloud of particles can form around the black hole Hod ; Herdeiro . The high spatial resolution, polarimetric imaging of supermassive black holes, like, for example, M87* or SgrA* by the Event Horizon Telescope can be used to prove the existence of ultralight bosonic particles Chen . The particles, existing around a rotating black hole due to the superradiance mechanism, concentrate in an accretion type disk. When linearly polarized photons are emitted from an accretion disk near the horizon, when traveling through the background cloud, their position angles oscillate due to the birefringent effect. The periodic change of the position angle can be tested observationally both spatially and temporally. The detection of such oscillations could give strong evidence for the existence of superradiance Chen . In this context the existence of the exact analytical solutions for the quasibound states could be very helpful for the understanding of the formation of the bosonic cloud, and of its properties.
The ultra light particles gravitating around black holes also have an effect on the gravitational waves emitted by the black holes in binary systems. Interesting enough, even a single gravitational - wave measurement can detect the ultra light bosons located around the gravitational wave source Hann . The observations of gravitational waves also show the existence of signals from the bosonic clouds formed by bound state of ultra light particles due to the spin-induced multipole moments and tidal Love numbers Bau .
Dyonic Kerr-Sen Black Hole have more variability as associated to their basic properties, as compared to the standard Schwarzschild or Kerr black holes, leading to a more complicated external dynamics. These richer properties do follow from the presence of the rotation, spin, and electric and magnetic charges, which also imply that the resulting black hole solution satisfies very complicated and strongly nonlinear field equations. As we have already seen, the effects associated with the spin and the charge degrees of freedom can lead to specific astrophysical effects and signatures, whose observational detection could open some new perspectives on the important relation between quantum and gravitational effects. In the present work, we have provided, via an exact solution of the Klein-Gordon equation in the Dyonic Kerr-Sen black hole background, some of the basic tools necessary for a detailed comparison of the predictions of the string inspired Einstein-Maxwell-dilaton-axion theory of gravity with the results of astrophysical observations. These results could also lead to a deeper understanding of the nature of the string theoretical effects, and of their physical relevance.
@@ Acknowledgments
We would like to thank Teephatai Bunyaratavej for pointing out an error in the earlier version of the paper. DS acknowledges this work is supported by the Second Century Fund (C2F), Chulalongkorn University, Thailand. P.B. is supported in part by National Research Council of Thailand (NRCT) and Chulalongkorn University under Grant N42A660500. This research has received funding support from the NSRF via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation [grant number B41G670027].
@@ Appendix A The Confluent Heun Functions and S-Homotopic Transformation
$y=e^{\pm i\sqrt{A_{5}}x}x^{\pm i\sqrt{A_{3}}}(x-1)^{\pm i\sqrt{A_{4}}}\left[A\operatorname{HeunC}\left(\alpha,\beta,\gamma,\delta,\eta,x\right)\right.\\ \left.+Bx^{-\beta}\operatorname{HeunC}\left(\alpha,-\beta,\gamma,\delta,\eta,x\right)\right],$ (182)